A snack food is made by mixing 2 lb of popcorn that costs $0.80 per pound with caramel that costs $2.90 per pound. How much caramel is needed to make a mixture that costs $2.20 per pound?
Answer by jorel555(1290)

2 lbs @ $0.80 per lb=1.60
2.9(n)+1.6=2.2(n+2)
2.9n+1.6=2.2n+4.4
.7n=2.8
n=2.8/.7=4 lbs of caramel!!!!!!!!